4. How many grams of Carbon dioxide are produced when 15g of o, are reacted with 30g of C,H16
Answer #1

Following is the - complete Answer -&- Explanation: for the given Question, in.....typed format....

\RightarrowAnswer:

Mass of carbon-di-oxide ( CO2 ), produced = 26.256 g ( grams )

\RightarrowExplanation:

Following is the complete Explanation, for the above Answer...

  • Given:
  1. ​​​​​​​Mass of n-heptane ( i.e. C7H16 ) =  30.0 g ( grams )
  2. Mass of oxygen ( O2 ) reacted = 15.0 g ( grams )
  • ​​​​​​​Step - 1:

​​​​​​​Following is the complete balanced reaction, between n-heptane ( C7H16 ), and oxygen ( O2 ):

\Rightarrowbalanced reaction: C7H16 + 11 O2 (g)   \rightleftharpoons 7 CO2 (g) + 8 H2O (l) -------------------- Equation - 1

\Rightarrow We get the following molar ratios:   ( i.e. from Equation - 1, above ...)

  1.   molar ratio:   C7H16 : O2 (g) =   1 : 11  
  2.   molar ratio:   C7H16 : CO2 (g) =   1 : 7  
  3.   molar ratio:    O2 : CO2 (g) = 11 : 7    
  • Step - 2:

​​​​​​​We know the following molar masses, of the given compounds:

  1. Molar mass of n-heptane ( C7H16 ) = 100.21 g/mol
  2. Molar mass of oxygen ( O2 ) = 16.0 g/mol
  3. Molar mass of carbon-di-oxide ( CO2 ) = 44.01 g/mol
  4. Molar mass of Water ( H2O ) = 18.0152 g/mol  
  • ​​​​​​​Step - 3:

​​​​​​​Therefore, using the above informtion, we will get the following:  i.e.  

  1. Number of moles of n-heptane ( C7H16 ) = ( mass ) / ( molar mass ) =  ( 30.0 g ) / ( 100.21 g/mol ) = 0.299 mol ( moles )
  2. Number of moles of oxygen ( O2 ) = (mass) / ( molar mass ) = ( 15.0 g ) / (16.0 g/mol ) = 0.9375 mol (moles )

\RightarrowTherefore, from the above discussed molar ratios, we are getting that , there are not enough number of moles of oxygen, reacting with the given number of moles , of n-heptane ( C7H16 ), because here Oxygen ( O2 ) , is serving the purpose of being the Limiting reactant.​​​​​​​

  • Step - 4:

​​​​​​​Therefore, we will use the molar ratio, we have obtained between: Oxygen ( O2 ), and Carbon-di-oxide ( CO2 ), in the above: Step - 1, to get our desired answer.

We know:

\Rightarrow molar ratio:    O2 : CO2 (g) = 11 : 7    

\Rightarrow  molar ratio: O2 : CO2 (g) = 1 : 7/11 = 1 : 0.6363 ( approx. )

\Rightarrowmolar ratio: O2 : CO2 (g) = 0.9375 : ( 0.6363 x 0.9375 ) =  0.9375 : 0.5966

  • Step - 5:

Therefore, from the above discussion, we are getting that: from 0.9375 moles of Oxygen ( O2 ), we will get : 0.5966 moles of Carbon -di-Oxide )( CO2 ) ....

Therefore:

\Rightarrow Mass of carbbon-di-oxide ( CO2 ), obtained/produced = ( 0.5966 mol CO2 ) x ( 44.01 g/mol CO2 )

=  26.256 g ( grams )

\Rightarrow Therefore, we are getting that 26.256 g ( grams ) of CO2 , will be produced in the given reaction...

------------------------------------------------------------------------------------------------------